### v1

# Left-Right Husimi Representation of Chaotic Resonance States: Invariance and Factorization

Abstract
For chaotic scattering systems we investigate the left-right Husimi representation, which combines left and right resonance states. We demonstrate that the left-right Husimi representation is invariant in the semiclassical limit under the corresponding closed classical dynamics, which we call quantum invariance. Furthermore, we show that it factorizes into a classical multifractal structure times universal quantum fluctuations. Numerical results for a dielectric cavity, the three-disk scattering system, and quantum maps confirm both the quantum invariance and the factorization.
I Introduction
Understanding the structure of resonance states in systems with a chaotic classical phase space is an essential goal in quantum chaos [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Much progress has been made on right resonance states, which are conjectured to be the product of a factor of classical origin giving a multifractal structure depending on the decay rate times a universal factor given by exponentially distributed quantum fluctuations [15, 16, 17, 18]. This factorization conjecture for right resonance states was studied in position space and in phase space using the Husimi representation. Recently, for the factor of classical origin a conditionally invariant measure with given decay rate was constructed based on a generalization of Ulam’s method [18], replacing previous approximate methods [19, 20]. In Ref. [18] it is applied to systems with full and partial escape [10], e.g. the three-disk scattering system (full escape) [21, 22, 23] and dielectric cavities (partial escape) [24].
There are also alternative representations of resonance states that are useful to reveal their properties. One such representation was introduced by Ermann, Carlo, and Saraceno [9], which we call left-right Husimi representation (LR-Husimi representation). This representation combines left and right resonance states and is naturally related to the spectral decomposition of the time-evolution operator of the open system. The LR-Husimi representation was introduced for systems with full escape to investigate the quantum analogue of the classical chaotic saddle. In this representation it was found that scarring on short periodic orbits is especially prominent [9]. A constructive semiclassical approach allows for describing resonance states in the LR-Husimi representation using scar functions on short periodic orbits [25, 26, 27, 28]. This was later extended to systems with partial escape [29, 30] and localization properties were studied [31, 32]. Also from a mathematical perspective there is an interest in the LR-Husimi representation [23, 33, 34]. For uniformly hyperbolic system with full escape it was shown that individual resonance states in LR-Husimi representation can be described by Ruelle resonances of zeta functions.
It is an open question whether the LR-Husimi representation can be factorized in a similar way as right resonance states, namely into a factor of classical origin times universal fluctuations. If so, what is the interpretation of the factor of classical origin? For systems with full escape it should be related to the invariant chaotic saddle, while for systems with partial escape the interpretation is open. The fluctuations were studied in Ref. [31] and significant deviations from an exponential distribution were found, which were interpreted as a localization effect. Instead, we expect that exponential fluctuations from left and right resonance states have to be combined. Are they correlated, what is the resulting distribution, and is it universal?
A further set of questions is related to a possible invariance of the LR-Husimi representation. For full escape this might be expected, as the LR-Husimi representation is supported by the invariant chaotic saddle. If so, is the LR-Husimi representation also invariant for systems with partial escape? Is it invariant under the classical dynamics with or without escape, i.e. the open or closed dynamics? Is this invariance exact or approximate and what happens in the semiclassical limit?
In this paper we show that the LR-Husimi representation of resonance states in classically chaotic systems with full and partial escape has two important properties: (i) invariance under the closed dynamics in the semiclassical limit, which we call quantum invariance; (ii) factorization with one factor of classical origin times universal quantum fluctuations. We derive the distribution of the fluctuations, based on the assumption of independent right and left fluctuations. We find that the factor of classical origin is obtained by combining the conditionally invariant measures corresponding to left and right resonance states. Numerical support for properties (i) and (ii) is given for a dielectric cavity (partial escape), the three-disk scattering system (full escape), and quantum maps with partial and full escape.
The paper is organized as follows: In Sec. II we review the definition of left and right resonance states and their LR-Husimi representation for open quantum maps and time-continuous scattering systems. In Sec. III we show that the LR-Husimi representation is invariant under the closed dynamics in the semiclassical limit. In Sec. IV we derive that the LR-Husimi representation can be factorized into a product of a factor of classical origin times universal fluctuations, which is numerically verified in Sec. V. Finally, we summarize our results and outline further directions in Sec. VI.
II Phase space representation of resonance states
We review the definition of left and right resonance states and their LR-Husimi representation for open quantum maps in Sec. II.1 and for time-continuous scattering systems in Sec. II.2. Specific example systems are introduced in Sec. II.3. We visualize their resonance states using the LR-Husimi representation in Sec. II.4 and discuss its basic properties.
II.1 Quantum maps
Quantum maps are well studied in quantum chaos and have a classical counterpart that is a time-discrete dynamical system [35]. For closed quantum maps the dynamics is given by a unitary time-evolution operator $U_{\text{cl}}$. We consider quantum maps with escape for which the time-evolution operator $U$ is non-unitary. If the system has a compact phase space, e.g. $\Gamma=[0,1)\times[0,1)$, $U$ can be represented by an $N\times N$ matrix, where $N$ is the dimension of the Hilbert space [36, 37]. The size $h$ of a Planck cell in phase space is then given by $h=1/N$. The time-evolution operator of the quantum map with escape is given by
$U=U_{\text{cl}}R_{\mathrm{refl}},$ (1)
where $R_{\mathrm{refl}}$ is a hermitian reflection operator that models the escape [5, 3, 38, 20].
If the non-unitary time-evolution operator $U$ is also non-normal, i.e. $U^{\dagger}U\neq UU^{\dagger}$, the left and right eigenvector of $U$ with the same eigenvalue are not identical. A right resonance state $|R_{n}\rangle$, $n=1,\ldots,N$, fulfills
$U|R_{n}\rangle=\lambda_{n}|R_{n}\rangle\,,$ (2)
with eigenvalue $\lambda_{n}$. It can be expressed as
$\lambda_{n}=\text{e}^{-\frac{\gamma_{n}}{2}}\text{e}^{-\text{i}\theta_{n}},$ (3)
where $\gamma_{n}$ is the decay rate and $\theta_{n}$ is a phase. The corresponding left resonance state $|L_{n}\rangle$ fulfills
$\langle L_{n}|U=\lambda_{n}\langle L_{n}|\,,$ (4)
with the same eigenvalue $\lambda_{n}$. Alternatively, left resonance states can be considered as right resonance states in the adjoint eigenvalue problem, $U^{\dagger}|L_{n}\rangle=\overline{\lambda_{n}}|L_{n}\rangle$ with $U^{\dagger}=R_{\mathrm{refl}}U_{\text{cl}}^{-1}$ according to Eq. 1 where the overbar denotes complex conjugation.
There is freedom in the choice of the normalization condition. One possibility is $\langle L_{n}|L_{n}\rangle=\langle R_{n}|R_{n}\rangle=1$. Another choice is $\langle L_{n}|R_{n}\rangle=1$ with $\langle L_{n}|L_{n}\rangle=\langle R_{n}|R_{n}\rangle$. The following equations are written such that they are valid for any normalization. Left and right resonance states form a biorthogonal system, i.e.
$\langle L_{m}|R_{n}\rangle=\delta_{mn}\langle L_{n}|R_{n}\rangle\,$ (5)
for $m,n=1,\ldots,N$ and fulfill the completeness relation
$\sum_{n=1}^{N}\frac{|R_{n}\rangle\langle L_{n}|}{\langle L_{n}|R_{n}\rangle}=\mathbb{1},$ (6)
if there is no degeneracy. The spectral decomposition of the non-unitary time-evolution operator $U$ is given by [9]
$U=\sum_{n=1}^{N}\lambda_{n}\frac{|R_{n}\rangle\langle L_{n}|}{\langle L_{n}|R_{n}\rangle}\,.$ (7)
As the structure of resonance states depends on their decay rate $\gamma_{n}$, we also give some information about their distribution. It depends on the type of escape of the system: (i) for full escape (some columns of $U$ are zero) most decay rates are larger than the natural decay rate $\gamma_{\mathrm{nat}}$ from classical dynamics [39, 40, 41, 42, 43, 44], (ii) for partial escape most decay rates are in the interval $[\gamma_{\mathrm{nat}},\gamma_{\mathrm{inv}}]$ with $\gamma_{\mathrm{inv}}$ being the inverse natural decay rate of the time-inverted classical dynamics [38, 45, 46, 20]. In the semiclassical limit they are conjectured to accumulate around the typical decay rate $\gamma_{\mathrm{typ}}$ [38].
For the comparison of quantum resonance states with the underlying classical dynamics one needs a phase-space representation of resonance states. Here, we use the Husimi representation [47, 48] which can be interpreted as a probability density on phase space. For a right resonance state $|R_{n}\rangle$ it is given by
$\mathcal{H}^{\mathrm{R}}_{n}(\boldsymbol{x})=\frac{|\langle\alpha(\boldsymbol{x})|R_{n}\rangle|^{2}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\,|\langle\alpha(\boldsymbol{x}^{\prime})|R_{n}\rangle|^{2}}\,,$ (8)
where $|\alpha(\boldsymbol{x})\rangle$ is a coherent state centered at a phase-space point $\boldsymbol{x}=(q,p)\in\Gamma$ with normalization $\langle\alpha(\boldsymbol{x})|\alpha(\boldsymbol{x})\rangle~{}=~{}1$. In Eq. 8 we use a normalization of the Husimi representation such that $\int_{\Gamma}\text{d}\boldsymbol{x}\,\mathcal{H}^{\mathrm{R}}_{n}(\boldsymbol{x})=1$. Equivalently, one can define the Husimi representation of the left resonance state $\langle L_{n}|$ as
$\mathcal{H}^{\mathrm{L}}_{n}(\boldsymbol{x})=\frac{|\langle L_{n}|\alpha(\boldsymbol{x})\rangle|^{2}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\,|\langle L_{n}|\alpha(\boldsymbol{x}^{\prime})\rangle|^{2}}\,.$ (9)
A combined representation from left and right resonance states uses the complex-valued left-right Husimi amplitude (LR-Husimi amplitude)
$h^{\mathrm{LR}}_{n}(\boldsymbol{x})=\frac{\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle}{\langle L_{n}|R_{n}\rangle}\,.$ (10)
From this the left-right Husimi representation (LR-Husimi representation) is obtained by
$\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})=\frac{|h^{\mathrm{LR}}_{n}(\boldsymbol{x})|}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\,|h^{\mathrm{LR}}_{n}(\boldsymbol{x}^{\prime})|}\,,$ (11)
which was introduced in Ref. [9]. We define the LR-Husimi representation to be normalized, $\int_{\Gamma}\text{d}\boldsymbol{x}\,\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})=1$, such that it can be interpreted as a probability density on phase space. The LR-Husimi representation can be expressed directly in terms of the Husimi representations of the left and right resonance states [9], Eqs. 8 and 9,
$\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})=\frac{\sqrt{\mathcal{H}^{\mathrm{L}}_{n}(\boldsymbol{x})\mathcal{H}^{\mathrm{R}}_{n}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\,\sqrt{\mathcal{H}^{\mathrm{L}}_{n}(\boldsymbol{x}^{\prime})\mathcal{H}^{\mathrm{R}}_{n}(\boldsymbol{x}^{\prime})}}\,,$ (12)
which will be used in the rest of this paper.
For quantum maps with time-reversal symmetry, specifically $U_{\text{cl}}=U_{\text{cl}}^{T}$ in position space, the left resonance states are given by $L_{n,q}=\sum_{q^{\prime}}(U_{\text{cl}})_{q,q^{\prime}}\overline{R_{n,q^{\prime}}}$ (this differs from Ref. [49], as we use an interchanged order of $U_{\text{cl}}$ and $R_{\mathrm{refl}}$ in Eq. 1). One can show that therefore the left Husimi representation is approximately related to the right Husimi representation by $\mathcal{H}^{\mathrm{L}}_{n}(q,p)\approx R_{\mathrm{refl}}(q,p)\mathcal{H}^{\mathrm{R}}_{n}(q,1-p)$, where $R_{\mathrm{refl}}(q,p)$ is the classical reflection at phase space point $(q,p)$. This approximation becomes better in the semiclassical limit, where the reflection operator $R_{\mathrm{refl}}$ applied to a coherent state is well approximated by the classical reflection function evaluated at the center of the coherent state. For the LR-Husimi representation, Eq. 12, this implies an approximate mirror symmetry in $p$, $\mathcal{H}^{\mathrm{LR}}(q,p)\approx\mathcal{H}^{\mathrm{LR}}(q,1-p)$, if additionally the reflection is symmetric, $R_{\mathrm{refl}}(q,p)=R_{\mathrm{refl}}(q,1-p)$.
II.2 Time-continuous scattering systems
Another class of systems which are especially of interest for experimental realizations are time-continuous scattering systems, in particular two-dimensional billiards with escape like dielectric cavities [24] or the three-disk scattering system [21, 50, 43, 23, 51]. In position space, resonance states are solutions of the Helmholtz equation fulfilling outgoing boundary conditions,
$\Delta\psi^{\mathrm{R}}(\boldsymbol{r})=-n^{2}(\boldsymbol{r})k^{2}\psi^{\mathrm{R}}(\boldsymbol{r})\,,$ (13)
for right resonance states $\psi^{\mathrm{R}}(\boldsymbol{r})$, where $\Delta$ is the Laplace operator in two dimensions, $n(\boldsymbol{r})$ is the local refractive index, and $k$ is the complex wavenumber. The imaginary part of $k$ is proportional to the dimensionless decay rate $\gamma$ [24, 17]. Left resonance states $\psi^{\mathrm{L}}(\boldsymbol{r})$ fulfill
$\psi^{\mathrm{L}}(\boldsymbol{r})\Delta=-n^{2}(\boldsymbol{r})k^{2}\psi^{\mathrm{L}}(\boldsymbol{r})$ (14)
with incoming boundary condition.
The LR-Husimi representation for time-continuous scattering systems is obtained by using the Husimi representation for right and left resonance states, see Eq. 12. For dielectric cavities the Husimi representation for right and left resonance states is defined in Ref. [52], specifically for right (left) resonance states we use the incoming (emanating) Husimi representation $\mathcal{H}_{n}^{\text{inc(em)}}$. For the three-disk scattering system the Husimi representation is defined in Ref. [23], where also the LR-Husimi representation is shown.
In analogy to the discussion of time-reversal symmetry for quantum maps, for the considered time-continuous scattering systems one has $\psi^{\mathrm{L}}_{n}(\boldsymbol{r})=\overline{\psi^{\mathrm{R}}_{n}(\boldsymbol{r})}$ [53, 23]. Therefore, the left Husimi representation is approximately $\mathcal{H}^{\mathrm{L}}_{n}(q,p)\approx R_{\mathrm{refl}}(q,p)\mathcal{H}^{\mathrm{R}}_{n}(q,-p)$. For the LR-Husimi representation, Eq. 12, this implies an approximate mirror symmetry in $p$, $\mathcal{H}^{\mathrm{LR}}(q,p)\approx\mathcal{H}^{\mathrm{LR}}(q,-p)$ using that the reflection $R_{\mathrm{refl}}(q,p)$ is symmetric in $p$ for billiard dynamics.
II.3 Example systems
We use four exemplary chaotic scattering systems throughout this paper [18]. These examples cover the cases of partial and full escape for a 2D billiard and a map in each case and have a fully chaotic phase space:
  • (a)
    Dielectric cavity (partial escape) with limaçon shape, deformation $\epsilon=0.6$, where it is practically fully chaotic, radius $R_{\mathrm{cav}}$, refractive index $n_{\mathrm{r}}=3.3$, and reflection law for a TM polarized mode [16]. The dimensionless decay rate $\gamma$ is related to the wavenumber $k$ by $\gamma=-2\,\mathrm{Im}\,kR_{\text{cav}}$.
  • (b)
    Three-disk scattering system (full escape) with disks of radius $a$ and center to center distance $2.1a$ [17]. The dimensionless decay rate is $\gamma=-2\,\mathrm{Im}\,ka$.
  • (c)
    Standard map at kicking strength $K=10$, where it is practically fully chaotic, with partial reflectivity $0.2$ in the interval $q\in[0.3,0.6]$ [20].
  • (d)
    Like (c) but with reflectivity 0 (full escape) [20].
All example systems have time-reversal symmetry (in the closed system) and the reflection has a mirror symmetry in $p$. For more details on the specific systems used here, see the supplemental material of Ref. [18].
[…]
II.4 Visualization of LR-Husimi representation
Figure 1: LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$ of individual resonance states closest to the specified decay rate $\gamma$ for (a) a dielectric cavity, (b) the three-disk scattering system, (c) the standard map with partial escape, and (d) the standard map with full escape. System parameters are given in Sec. II.3. In this and all following figures the average intensity is scaled to one (in (a) and (c) on the full phase space and in (b) and (d) on the chaotic saddle). Intensities greater than the maximal value of the colorbar are shown with darkest color.
Fig. 1 shows the LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$ for resonance states of all example systems for various decay rates. Due to the above discussed symmetry in all example systems, the LR-Husimi representation is approximately mirror symmetric in $p$.
For systems with partial escape, Fig. 1(a, c), the LR-Husimi representation is supported by the entire phase space. In contrast, for systems with full escape, Fig. 1(b, d), it is localized on the chaotic saddle. This can be understood as follows: The Husimi representation of left (right) resonance states is supported by the forward (backward) trapped set [3, 6]. The LR-Husimi representation is a product of both Husimi representations, Eq. 12, and hence is supported by the intersection of both sets, which is the chaotic saddle (sometimes termed repeller).
Furthermore, for systems with full escape, Fig. 1(b, d), the LR-Husimi representation is almost independent of the decay rate $\gamma$ compared to the strong dependence on $\gamma$ for partial escape, Fig. 1(a, c). More precisely, for all decay rates it is constant on the chaotic saddle up to fluctuations (with a $\sqrt{1-p^{2}}$ dependence for the three-disk scattering system due to the chosen definition of the boundary Husimi representation [54]).
III Quantum invariance
Figure 2: Visual comparison of the LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$ according to left-hand side (left) and right-hand side (right) of Eq. 16. The integration is done over regions $A$ from (a) a fine $1600\times 1600$ grid showing deviations from invariance due to stretching along the unstable direction, see zoom by factor 16, and (b) a coarse $50\times 50$ grid showing less deviations from invariance (even though we choose a more sensitive range for the colorbar). The presented individual resonance state is closest to $\gamma=0.03$ for a dielectric cavity, see Fig. 1(a, second panel).
Figure 3: Semiclassical convergence of left-hand side to right-hand side of Eq. 16 for a fixed size of regions $A$ of a $50\times 50$ grid demonstrating quantum invariance. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ evaluated on this grid approaching zero in the semiclassical limit. (a-d) Four example systems with resonance states near five decay rates each, as in Fig. 1. Symbols as in Fig. 5.
We demonstrate that resonance states in the LR-Husimi representation are invariant under the closed dynamics in the semiclassical limit. We call this quantum invariance in analogy to quantum ergodicity, as explained below. First we present visual and quantitative numerical evidence, later we give analytical support.
Invariance of some phase-space distribution $\rho(\boldsymbol{x})$ under a map $T$ is defined by
$\int_{A}\text{d}\boldsymbol{x}\,\rho(\boldsymbol{x})=\int_{T^{-1}(A)}\text{d}\boldsymbol{x}\,\rho(\boldsymbol{x})\,,$ (15)
for any phase-space region $A$, where $T^{-1}(A)$ is the preimage of $A$. In the present context the classical map $T$ corresponds to the unitary time-evolution operator $U_{\text{cl}}$ of the closed quantum map. Note that $T$ also can be the boundary map of a billiard system. The time it takes from one to the next reflection in the billiard has no relevance for invariance.
When specifying such an invariance for the LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$,
$\int_{A}\text{d}\boldsymbol{x}\,\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})\approx\int_{T^{-1}(A)}\text{d}\boldsymbol{x}\,\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})\,,$ (16)
it is at best approximately true. This is obvious for regions $A$ smaller than or of the same size as a coherent state, i.e. the scale $h$ of a Planck cell: The Husimi representation is smooth on this scale and it has local maxima and zeros with the same elliptic shape as the coherent states used in the representation. Under the hyperbolic dynamics this shape becomes stretched and leads to deviations from invariance in Eq. 16. This is visualized for a resonance state of the dielectric cavity in Fig. 2, comparing the left-hand side and right-hand side of Eq. 16 shown on the left and the right, respectively. In Fig. 2(a) we evaluate Eq. 16 for regions $A$ of a fine $1600\times 1600$ grid in phase space, i.e. $A\ll h$. On the right one observes a stretching along the unstable direction, best seen in the zoom region. In this case one indeed finds deviations from invariance. In contrast, in Fig. 2(b) we use regions $A$ of a coarse $50\times 50$ grid. For these larger regions $A\gg h$ the stretched shape on the right is no longer resolved. The left-hand side and right-hand side of Eq. 16 appear to be much closer.
This is analyzed quantitatively for a fixed size of the regions $A$ when going to the semiclassical limit, $h\rightarrow 0$, choosing resonance states near a fixed decay rate $\gamma$, see Fig. 3. Specifically, we use regions $A$ from a $50\times 50$ grid and quantify the difference between the left-hand side and right-hand side of Eq. 16, as visualized in Fig. 2(b), using the Jensen-Shannon divergence [55]. In the semiclassical limit, i.e. larger wavenumber $k$ or larger matrix size $N$, it decays to zero for all example systems and decay rates.
We thus define quantum invariance as a property of resonance states in their LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$: In the semiclassical limit $h\rightarrow 0$ and for a phase-space region $A$ with a fixed size, there is invariance under the closed dynamics, i.e. the left-hand side and right-hand side of Eq. 16 approach each other. We use the term quantum invariance in analogy to the term quantum ergodicity, where a property from dynamical systems is attributed to a quantum system under certain qualifications in the semiclassical limit.
We mention as an aside that we attribute the different magnitude of the Jensen-Shannon divergence in Fig. 3 to the different magnitude of the Lyapunov exponent in the various systems. In particular, the Lyapunov exponent is the smallest for the limaçon shaped dielectric cavity, leading to the least stretching and thus to the least deviations from invariance, i.e. the smallest value of the Jensen-Shannon divergence.
In the remainder of this section we demonstrate analytically that the LR-Husimi representation is approximately invariant under the closed dynamics, supporting quantum invariance. We discuss the case of a quantum map from Sec. II.1 with a time-evolution operator according to Eq. 1, $U=U_{\text{cl}}R_{\mathrm{refl}}$, where $U_{\text{cl}}$ is the unitary time-evolution operator of the closed map. In the numerator of the definition of the LR-Husimi amplitude, Eq. 10, (i) we first insert the identity $I=U_{\text{cl}}U_{\text{cl}}^{-1}$ twice, leading to
$\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle$ =$\stackrel{{\scriptstyle(\text{i})}}{{=}}\langle\alpha(\boldsymbol{x})|U_{\text{cl}}U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle$
$\stackrel{{\scriptstyle(\text{ii})}}{{\approx}}\langle\alpha(T^{-1}(\boldsymbol{x}))|U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
=$\stackrel{{\scriptstyle(\text{iii})}}{{=}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
$\stackrel{{\scriptstyle(\text{iv})}}{{\approx}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{n}\rangle\langle L_{n}|\alpha(T^{-1}(\boldsymbol{x}))\rangle\,.$ (17)
In step (ii) we use the approximate identity
$U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle\approx|\alpha(T^{-1}(\boldsymbol{x}))\rangle\,,$ (18)
that the backward time-evolved coherent state is (up to squeezing) the coherent state at the backward time-evolved phase-space position. In step (iii) we use the identity
$U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}=R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}\,,$ (19)
which is based on the definition of right and left eigenstates (product of applying $U_{\text{cl}}^{-1}$ from the left to Eq. 2 and $R_{\mathrm{refl}}^{-1}$ from the right to Eq. 4; for systems with full escape, where $R_{\mathrm{refl}}^{-1}$ is not defined, one has to suitably restrict the argument to the chaotic saddle). In the last step (iv) of Eq. 17 we use that the reflection function is approximately constant on the scale of the coherent state. Then for $\boldsymbol{y}=T^{-1}(\boldsymbol{x})$ the term $R_{\mathrm{refl}}^{-1}|\alpha(\boldsymbol{y})\rangle$ multiplies the coherent state with a factor that is the inverse of the factor from $\langle\alpha(\boldsymbol{y}))|R_{\mathrm{refl}}$ such that they cancel.
From Eq. 17 follows the approximate pointwise invariance of the smooth LR-Husimi amplitude, Eq. 10,
$h^{\mathrm{LR}}_{n}(\boldsymbol{x})\approx h^{\mathrm{LR}}_{n}(T^{-1}(\boldsymbol{x}))\,.$ (20)
By integrating over some phase-space region $A$, we find the approximate invariance,
$\int_{A}\text{d}\boldsymbol{x}\,h^{\mathrm{LR}}_{n}(\boldsymbol{x})\approx\int_{A}\text{d}\boldsymbol{x}\,h^{\mathrm{LR}}_{n}(T^{-1}(\boldsymbol{x}))=\int_{T^{-1}(A)}\text{d}\boldsymbol{y}\,h^{\mathrm{LR}}_{n}(\boldsymbol{y})\,.$
Similarly, one finds for the LR-Husimi representation, using $\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})\propto|h^{\mathrm{LR}}_{n}(\boldsymbol{x})|$ according to Eq. 12, an approximate invariance corresponding to Eq. 16. In order to demonstrate quantum invariance one would have to show that in the semiclassical limit, when the size of the coherent state is arbitrarily small compared to the region $A$, the relative error of the used approximations goes to zero when integrating over $A$.
IV Factorization
In this section we show that the LR-Husimi representation of resonance states can be factorized into a product of a density from classical dynamics and universal fluctuations. To this end we first review the factorization for right and left resonance states in Sec. IV.1. Based on this we derive the factorization for the LR-Husimi in Sec. IV.2 and analyze both factors in detail in Secs. IV.3 and IV.4.
IV.1 Factorization conjecture of resonance states
According to a recently proposed factorization conjecture [15, 16, 17], right resonance states in chaotic scattering are a product of (i) a conditionally invariant measure from classical dynamics which by smoothing on the scale of a Planck cell gives a smooth density $\varrho_{\gamma}^{\mathrm{R}}$ depending on the decay rate $\gamma$ and (ii) universal exponentially distributed fluctuations with mean one. This factorization holds for the intensity $|\langle\varphi|R\rangle|^{2}$ of right resonance states and is independent of the chosen (possibly over-complete) basis $|\varphi\rangle$ [15]. For example, using coherent states $|\varphi\rangle=|\alpha(\boldsymbol{x})\rangle$, the Husimi representation of a right resonance state with decay rate $\gamma$ is given by the product
$\mathcal{H}^{\mathrm{R}}(\boldsymbol{x})=\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})\cdot\eta^{\mathrm{R}}(\boldsymbol{x})\,.$ (21)
The fluctuations $\eta^{\mathrm{R}}$, statistically independent from $\varrho_{\gamma}^{\mathrm{R}}$, follow an exponential distribution $P(\eta^{\mathrm{R}})=\exp(-\eta^{\mathrm{R}})\,$.
Since left resonance states are right resonance states of a modified system they factorize as well,
$\mathcal{H}^{\mathrm{L}}(\boldsymbol{x})=\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\cdot\eta^{\mathrm{L}}(\boldsymbol{x})\,,$ (22)
with fluctuations $\eta^{\mathrm{L}}$, statistically independent from $\varrho_{\gamma}^{\mathrm{L}}$, with the same exponential distribution $P(\eta^{\mathrm{L}})=\exp(-\eta^{\mathrm{L}})\,$.
IV.2 Factorization of LR-Husimi representation
In the following, we derive such a factorization also for the LR-Husimi representation based on the factorization conjecture for left and right resonance states. To this end we start with Eq. 12 and insert Eqs. 21 and 22 to find
$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x})\eta^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}$
We now use from the factorization conjecture above that the right fluctuations $\eta^{\mathrm{R}}(\boldsymbol{x})$ are statistically independent of their classical density $\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})$. The same is true for the left fluctuations $\eta^{\mathrm{L}}(\boldsymbol{x})$ and their classical density $\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})$. Hence, the factor $\sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}$ in the denominator can be replaced by its expectation value. This leads to the factorization for the LR-Husimi representation
$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})\cdot\eta^{\mathrm{LR}}(\boldsymbol{x})\,,$ (24)
where the first factor is a smooth density,
$\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}\,.$ (25)
It is given by the square root of the product of the smooth densities $\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})$ and $\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})$ that describe the left and right resonance states, respectively. Below in Sec. IV.4, we will show that there is a measure $\mu_{\gamma}^{\mathrm{LR}}$ constructed from classical dynamics which by smoothing on the scale of a Planck cell gives $\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})$.
We call the second factor in Eq. 24 the LR-fluctuations, defined by
$\eta^{\mathrm{LR}}=\frac{\sqrt{\eta^{\mathrm{L}}\eta^{\mathrm{R}}}}{\text{E}\left[\sqrt{\eta^{\mathrm{L}}\eta^{\mathrm{R}}}\right]}\,.$ (26)
It is given by the square root of the product of two exponentially distributed random variables $\eta^{\mathrm{L}}$ and $\eta^{\mathrm{R}}$ with mean one. From their definition it follows directly that their expectation value is one, $\text{E}\left[\eta^{\mathrm{LR}}\right]=1$. Averaging Eq. 24 over resonance states with similar decay rates $\gamma$ thus relates the smooth density $\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})$ to the averaged LR-Husimi representation,
$\langle\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})\rangle_{\gamma}=\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})\,.$ (27)
In Fig. 4 we illustrate the factorization of the LR-Husimi representation for a resonance state for all example system, with the numerical approach explained in Sec. V. The first factor (middle column) will be compared to a measure $\mu_{\gamma}^{\mathrm{LR}}$ in Sec. V.3. The LR-fluctuations (right column) appear to be the same in the entire phase space for partial escape, Fig. 4(a, c), and on the chaotic saddle for full escape, Fig. 4(b, d). Given that the systems are very different, this suggests that the LR-fluctuations are universal. We derive their distribution in Sec. IV.3 and numerically support the universality in Sec. V.2.
Figure 4: Factorization of LR-Husimi representation $\mathcal{H}^{\mathrm{LR}}$ (left) according to Eq. 32 into the average $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ of resonance states close to decay rate $\gamma$ (middle) times LR-fluctuations $\eta^{\mathrm{LR}}$, Eq. 33, (right) for (a) dielectric cavity at $\gamma=0.053$, (b) three-disk scattering system at $\gamma=1.0$, (c) standard map with partial escape at $\gamma=0.055$, and (d) standard map with full escape at $\gamma=0.5$.
IV.3 Distribution of LR-fluctuations
We now derive the distribution of the LR-fluctuations, Eq. 26. They are defined by the square root of the product of two exponentially distributed random variables $\eta^{\mathrm{L}}$ and $\eta^{\mathrm{R}}$ with mean one and are normalized such that $\text{E}\left[\eta^{\mathrm{LR}}\right]=1$.
In the following we make the assumption of statistical independence of the left and right fluctuations, $\eta^{\mathrm{L}}$ and $\eta^{\mathrm{R}}$. Note that for systems with time-reversal symmetry (in the closed system) this is not applicable in position space, where left and right resonance states are complex conjugates of each other and therefore their fluctuations are maximally correlated.
First, we evaluate the denominator in Eq. 26,
$\text{E}\left[\sqrt{\eta^{\mathrm{L}}\eta^{\mathrm{R}}}\right]=\left(\int_{0}^{\infty}\text{d}\eta\,P(\eta)\,\sqrt{\eta}\right)^{2}=\frac{\pi}{4}\,,$ (28)
where $P(\eta)=\exp(-\eta)$. The probability distribution $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})$ is then given by
$P^{\mathrm{LR}}(\eta^{\mathrm{LR}})=\int_{0}^{\infty}\text{d}\eta^{\mathrm{L}}\int_{0}^{\infty}\text{d}\eta^{\mathrm{R}}\,P(\eta^{\mathrm{L}})\,P(\eta^{\mathrm{R}})\;\;\delta\left(\eta^{\mathrm{LR}}-\frac{4}{\pi}\sqrt{\eta^{\mathrm{L}}\eta^{\mathrm{R}}}\right)\,.$ (29)
For one integration the $\delta$-function can be exploited, while the other gives an integral representation of $K_{0}$, the 00-th order modified Bessel function of the second kind [56], leading to the universal distribution of LR-fluctuations,
$P^{\mathrm{LR}}(\eta^{\mathrm{LR}})=\frac{\pi^{2}}{4}\,\eta^{\mathrm{LR}}\,K_{0}\left(\frac{\pi}{2}\eta^{\mathrm{LR}}\right)\,.$ (30)
Figure 5: Convergence of coarse-grained LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ for increasingly fine partitions with $n$ cells. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ (evaluated on a $50\times 50$ grid) between measures from partitions $n$ and $n/4$ approaching zero. (a) Dielectric cavity for $\gamma\in\{0.011\,(\gamma_{\mathrm{nat}}),0.030,0.053\,(\gamma_{\mathrm{typ}}),0.090,0.122\,(\gamma_{\mathrm{inv}})\}$. (b) Three-disk scattering system for $\gamma\in\{0.436\,(\gamma_{\mathrm{nat}}),0.6,1.0,1.4,1.8\}$. (c) Standard map with partial escape for $\gamma\in\{0.22\,(\gamma_{\mathrm{nat}}),0.35,0.55,0.75,0.88\,(\gamma_{\mathrm{inv}})\}$. (d) Standard map with full escape for $\gamma\in\{0.25\,(\gamma_{\mathrm{nat}}),0.35,0.5,0.75,1.0\}$. Symbols $\circ$, $+$, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.
IV.4 Construction of LR-measure
Next, we propose the construction of a measure $\mu_{\gamma}^{\mathrm{LR}}$, which we call LR-measure. This multifractal measure, after smoothing on the scale of a Planck cell, gives the smooth density $\varrho_{\gamma}^{\mathrm{LR}}(\boldsymbol{x})$Eq. 25. In Sec. V.3 the LR-measure is compared to the first factor of the factorization in Fig. 4.
In Ref. [18] we proposed the construction of a conditionally invariant measure $\mu_{\gamma}^{\mathrm{R}}$ that describes the average of right resonance states with decay rate $\gamma$ in the semiclassical limit. Its construction generalizes Ulam’s method [57] for finding the leading eigenvector of a matrix approximation of the Perron-Frobenius operator [51]. This is done by allowing more general matrix approximations of the operator [18]. From the infinite set of matrix approximations one has to select the one that fulfills an optimization criterion [18]. The conditionally invariant measure $\mu_{\gamma}^{\mathrm{R}}$ is then given by the leading eigenvector of the selected matrix approximation. Numerically, for a partition of phase space into $n$ disjoint cells $\{A_{1},\dots,A_{n}\}$ one determines the coarse-grained conditionally invariant measure $\mu_{\gamma,i}^{\mathrm{R}}=\mu_{\gamma}^{\mathrm{R}}(A_{i})$. In the limit of increasingly finer partitions of phase space, $n\rightarrow\infty$, the right measures $\mu_{\gamma}^{\mathrm{R}}$ were found to converge [18].
The proposed method can also be used to find a conditionally invariant measure $\mu_{\gamma}^{\mathrm{L}}$ that describes left resonance states. Since left resonance states are right resonance states of $U^{\dagger}$, see discussion after Eq. 4, one uses the time-reversed closed dynamics and applies the reflection afterwards. For systems with time-reversal symmetry (in the closed system), the left measure $\mu_{\gamma}^{\mathrm{L}}$ can be directly obtained from the right measure in phase space by mirroring in $p$ and subsequently applying the reflection.
We define the coarse-grained LR-measure $\mu_{\gamma,i}^{\mathrm{LR}}=\mu_{\gamma}^{\mathrm{LR}}(A_{i})$ for each cell $A_{i}$ by box-wise multiplication of the left and right measures and normalization,
$\mu_{\gamma,i}^{\mathrm{LR}}=\frac{\sqrt{\mu_{\gamma,i}^{\mathrm{L}}\ \mu_{\gamma,i}^{\mathrm{R}}}}{\sum_{i^{\prime}=1}^{n}\sqrt{\mu_{\gamma,i^{\prime}}^{\mathrm{L}}\ \mu_{\gamma,i^{\prime}}^{\mathrm{R}}}}\,.$ (31)
The square root is taken in accordance to Eq. 25. Note that we use boxes of equal size, otherwise one would have to suitably adopt the normalization.
In the limit of increasingly finer partitions of phase space, $n\rightarrow\infty$, we find convergence for the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$, see Fig. 5. We numerically demonstrate the convergence by comparing the coarse-grained measures for $n$ and for $n/4$ cells using the Jensen-Shannon divergence [55] (its square root being a metric). Figure 5 shows that with increasing $n$ the Jensen-Shannon divergence converges to zero. Here, the finest phase-space partition has $n=3200^{2}$ cells. Furthermore, the ratio of two consecutive Jensen-Shannon divergences is bounded from above by a value smaller than one. If this continues to hold for increasing $n$, the coarse-grained measures are a contractive sequence, therefore a Cauchy sequence and thus converge in the limit $n\rightarrow\infty$. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma>\gamma_{\text{nat}}$ (full escape) for all example systems.
Note that for the natural decay rate $\gamma_{\mathrm{nat}}$ (or the inverse natural decay rate $\gamma_{\mathrm{inv}}$) the construction of the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ is also possible by a simpler method: One can generate the right natural measure by a long-term time evolution of many phase-space points and their intensities [3, 4, 10]. Equivalently, this can be done for the left natural measure by time-reversed dynamics. Correspondingly, one finds the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ for $\gamma_{\mathrm{nat}}$ by iterating phase-space points forward and backward in time, multiplying the intensities, and taking the square root. This was previously done in Ref. [58, Fig. 8] (without taking the square root) in order to extend the concept of the chaotic saddle to the case of partial escape.
V Numerical Verification of Factorization
In this section we numerically verify the factorization of the LR-Husimi representation. First, we explain in Sec. V.1 how we numerically factorize the LR-Husimi representation into an average times fluctuations. In Sec. V.2 we compare the LR-fluctuations to the theoretically predicted distribution and verify their universality in the semiclassical limit. Furthermore, we demonstrate in Sec. V.3 that the averaged LR-Husimi representation is well described by the LR-measure and find convergence in the semiclassical limit.
V.1 Factorization
For the numerical verification of the factorization of the LR-Husimi representation, Eq. 24, we use for the first factor the averaged LR-Husimi representation, Eq. 27, giving
$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\langle\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})\rangle_{\gamma}\cdot\eta^{\mathrm{LR}}(\boldsymbol{x})\,.$ (32)
From this we compute the LR-fluctuations by
$\eta^{\mathrm{LR}}(\boldsymbol{x})=\frac{\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})}{\langle\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})\rangle_{\gamma}}\,.$ (33)
Note that for systems with full escape the LR-fluctuations $\eta^{\mathrm{LR}}$ should be evaluated on the chaotic saddle only, as the denominator tends to zero elsewhere.
Figure 4 shows the factorization of one resonance state in LR-Husimi representation for all example systems. For averaging we use 500 resonance states which are close to the specified decay rate $\gamma$.
Let us mention that the original way of defining the LR-Husimi representation [9] is without the normalization of Eq. 11, $\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})=\left|h^{\mathrm{LR}}_{n}(\boldsymbol{x})\right|$, such that $\int_{\Gamma}\text{d}\boldsymbol{x}\,\mathcal{H}^{\mathrm{LR}}_{n}(\boldsymbol{x})$ varies for each resonance state. This might affect the averaged LR-Husimi representation and thus the LR-fluctuations. Numerically, however, we find no significant consequences.
V.2 LR-fluctuations
Figure 6: Visual comparison of the histogram of LR-fluctuations $\eta^{\mathrm{LR}}$, Eq. 33, to the predicted probability distribution $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})$, Eq. 30 (black), on linear scale (top) and logarithmic scale (bottom). The resonance states from Fig. 1 with five different decay rates (colored) are used for (a) dielectric cavity, (b) three-disk scattering system, (c) standard map with partial escape, and (d) standard map with full escape.
Figure 7: Convergence of the distribution of LR-fluctuations $\eta^{\mathrm{LR}}$, Eq. 33, to theoretically predicted distribution $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})$, Eq. 30, for systems and decay rates of Fig. 6. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ (evaluated on 100 bins over the interval $[0,10]$) approaching zero in the semiclassical limit.
We now give numerical evidence that the LR-fluctuations $\eta^{\mathrm{LR}}$, Eq. 33, follow the theoretically predicted distribution $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})$Eq. 30. This is shown in Fig. 6 for all example systems and demonstrates the universality of the LR-fluctuations.
In the semiclassical limit the distribution of the LR-fluctuations $\eta^{\mathrm{LR}}$ should converge to the predicted distribution $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})$ from Eq. 30. We verify this using the Jensen-Shannon divergence [55] as a quantification of the distance of the distributions. Figure 7 shows that going further towards the semiclassical limit the Jensen-Shannon divergence converges to zero. It is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma>\gamma_{\text{nat}}$ (full escape) for all example systems. A posteriori, this justifies our assumption of statistical independence of the left and right fluctuations $\eta^{\mathrm{L}}$ and $\eta^{\mathrm{R}}$ in Sec. IV.3. In fact, we numerically checked that the Pearson correlation coefficient between the left and right fluctuations $\eta^{\mathrm{L}}$ and $\eta^{\mathrm{R}}$ decays in the semiclassical limit (not shown).
These results on the distribution of LR-fluctuations explain some of the findings of Ref. [31]. In particular, the comparison of the LR-fluctuations to an exponential distribution shows strong deviations, as shown in Ref. [31, Fig. 1]. This observation is consistent with the above prediction that the LR-fluctuations are distributed according to Eq. 30. Furthermore, in Ref. [31, Fig. 2] a localization measure shows differences between the LR-fluctuations and the fluctuations of the Husimi representation of right resonance states. We attribute this to the different underlying universal distributions, namely Eq. 30 compared to the exponential distribution.
V.3 Averaged LR-Husimi representation
Figure 8: Visual comparison of the averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ (upper half) to the proposed LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ smoothed on the scale of a Planck cell (lower half) for five different decay rates. For the average we use 500 resonance states close to the given decay rate $\gamma$. This is shown for (a) dielectric cavity, (b) three-disk scattering system, (c) standard map with partial escape, and (d) standard map with full escape.
Figure 9: Convergence of averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ to LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ for systems and decay rates of Fig. 8. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ (evaluated on a $50\times 50$ grid) approaching zero in the semiclassical limit.
In the following we show that the averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ is described by the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$. To this end we first give a visual comparison for all example systems and various decay rates $\gamma$ in Fig. 8. Here the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ is computed on a $3200\times 3200$ grid ($n\approx 10^{7}$ partition cells). For visual comparison we smooth it on the scale corresponding to the Planck cell $h$ of the LR-Husimi representation. We see an overall good agreement between the averaged LR-Husimi representation and the LR-measure for all cases. Small differences appear only on the scale of a Planck cell, see e.g. $\gamma_{\mathrm{typ}}$ (middle) in Fig. 8(a) and the largest decay rate in Fig. 8(b).
An even better visual comparison can be achieved by first smoothing the individual left and right measure before their multiplication in Eq. 31, see Appendix A. However, this is not necessary in order to find agreement in the semiclassical limit, see below.
For systems with full escape we observe that the LR-measure does not depend on $\gamma$ and that it is uniform on the chaotic saddle. We explain this using the fact that the dynamical system confined to the chaotic saddle is closed and ergodic implying a uniform distribution of the LR-measure. Note that for the three-disk scattering system the LR-measure is multiplied by $\sqrt{1-p^{2}}$ for the visual comparison as this factor is a property of the boundary Husimi representation [54].
In the semiclassical limit the averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ should converge to the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$. We verify this using the Jensen-Shannon divergence as a quantification of the distance between $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ and $\mu_{\gamma}^{\mathrm{LR}}$. Figure 9 shows that going further towards the semiclassical limit the Jensen-Shannon divergence converges to zero. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma>\gamma_{\text{nat}}$ (full escape) for all example systems.
VI Summary and Outlook
In summary, we demonstrate the quantum invariance of the LR-Husimi representation and show that it can be factorized into a product of a smooth classical density $\varrho_{\gamma}^{\mathrm{LR}}$ and universal LR-fluctuations $\eta^{\mathrm{LR}}$. Quantum invariance is introduced as a property of resonance states in their LR-Husimi representation and states that in the semiclassical limit $h\to 0$ and for a phase-space region with fixed size it is invariant under the closed dynamics. In the factorization of the LR-Husimi representation the first factor $\varrho_{\gamma}^{\mathrm{LR}}$ is a classical density, smooth on the scale of a Planck cell, determined from the conditionally invariant measures of the left and right resonance states. Its structure strongly depends on the specific system. For systems with partial escape there is also a strong dependence on the decay rate. In contrast, for systems with full escape, $\varrho_{\gamma}^{\mathrm{LR}}$ is uniform on the chaotic saddle and is independent of $\gamma$. The second factor $\eta^{\mathrm{LR}}$ describes fluctuations around this average and follows a universal distribution, Eq. 30, that is independent of the system and the decay rate. The derivation of this distribution is based on the assumption of statistical independence of the left and right fluctuations. We justify this assumption a posteriori by the numerical verification of the distribution.
We expect that the coarse-grained LR-measure fulfills invariance under the closed dynamics similar to the quantum invariance of the LR-Husimi representation. Here, the finite-sized phase-space partition of Ulam’s method replaces the finite-sized Planck cell. In the limit of an infinitely fine partition this should lead to an invariant measure $\mu_{\gamma}^{\mathrm{LR}}$ under the closed dynamics, which could be tested numerically. If so, then for systems with partial escape one finds a one-parameter family of invariant measures depending on the decay rate $\gamma$. It is an open question whether or not these are related to the invariant measures (equilibrium states) of the thermodynamic formalism in ergodic theory, which depend on the inverse temperature $\beta$ [59, 60, 61].
Furthermore, it is an interesting open problem whether the presented results on quantum invariance for fully chaotic system generalize to systems with a mixed (regular and chaotic) phase space. The analytical arguments in Sec. III do not depend on the type of dynamics and thus seem to suggest that quantum invariance holds more generally for resonance states with a well defined semiclassical limit.
Acknowledgements.
We are grateful for discussions with A. Bäcker, E.-M. Graefe, and D. Savin. RK thanks M. M. Castro and G. Froyland for illuminating discussions on invariant measures and the possible relation of this work to equilibrium states of ergodic theory. We thank L. Müller for contributions in the early stage of the project [62]. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 262765445. The authors gratefully acknowledge the computing time on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR.
Figure 10: Visual comparison of the averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ (upper half, same as in Fig. 8) to an alternative LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ with smoothing on the scale of a Planck cell before multiplication (lower half) for five different decay rates. This is shown for (a) dielectric cavity, (b) three-disk scattering system, (c) standard map with partial escape, and (d) standard map with full escape.
Appendix A Alternative smoothing of LR-measure
The comparison of the averaged LR-Husimi representation $\langle\mathcal{H}^{\mathrm{LR}}\rangle_{\gamma}$ to the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ in Fig. 9 convergences in the semiclassical limit, but the visual comparison for finite wavelengths in Fig. 8 shows deviations on the scale of a Planck cell. This is improved by smoothing the individual left and right measures $\mu_{\gamma}^{\mathrm{L}}$ and $\mu_{\gamma}^{\mathrm{R}}$ on the scale of a Planck cell before multiplying them according to Eq. 31, see Fig. 10. We motivate this by the fact that the individual left and right Husimi representations are also smooth on the scale of a Planck cell before being multiplied in Eq. 12 to obtain the LR-Husimi representation.
At finite wavelengths this alternative construction of the LR-measure is better able to reproduce the intensity maxima in the LR-Husimi representation, e.g. compare Figs. 10(a) and 8(a) at $\gamma_{\mathrm{typ}}$ (middle) or Figs. 10(b) and 8(b) at the largest decay rate. Additionally, for full escape the smallest scale of the multifractal structure is better reproduced in Fig. 10(b, d) than in Fig. 8(b, d) for all decay rates.
However, we also observe a drawback of this alternative construction. For the three-disk scattering system, Fig. 10(b), one observes a dependence of the measure $\mu_{\gamma}^{\mathrm{LR}}$ on the $s$-direction for $p\approx 0$ for larger decay rates. This does not occur for the averaged LR-Husimi representation. Additionally, it has the disadvantage that the construction of the LR-measure depends on the size of the Planck cell. This is why we prefer the construction of the LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ in the main text, Eq. 31, with smoothing just for visualization purposes.

### v2

@@
 We consider quantum maps with escape for which the time-evolution operator $U$ is non-unitary.
-If the system has a compact phase space, e.g.
-$\Gamma=[0,1)\times[0,1)$, $U$ can be represented by an $N\times N$ matrix, where $N$ is the dimension of the Hilbert space [36, 37].
+If the system has a two-dimensional compact phase space, e.g.
+the torus $\Gamma=[0,1)\times[0,1)$, $U$ can be represented by an $N\times N$ matrix, where $N$ is the dimension of the Hilbert space [36, 37].
 The size $h$ of a Planck cell in phase space is then given by $h=1/N$.
+Note that this can be straightforwardly generalized to a higher-dimensional phase space.
 The time-evolution operator of the quantum map with escape is given by
@@
 As the structure of resonance states depends on their decay rate $\gamma_{n}$, we also give some information about their distribution.
-It depends on the type of escape of the system: (i) for full escape (some columns of $U$ are zero) most decay rates are larger than the natural decay rate $\gamma_{\mathrm{nat}}$ from classical dynamics [39, 40, 41, 42, 43, 44], (ii) for partial escape most decay rates are in the interval $[\gamma_{\mathrm{nat}},\gamma_{\mathrm{inv}}]$ with $\gamma_{\mathrm{inv}}$ being the inverse natural decay rate of the time-inverted classical dynamics [38, 45, 46, 20].
+It depends on the type of escape of the system: (i) for full escape (some columns of $U$ are zero) most decay rates are larger than the natural decay rate $\gamma_{\mathrm{nat}}$ from classical dynamics [39, 40, 41, 42, 43, 44], (ii) for partial escape most decay rates are in the interval $\gamma_{\mathrm{nat}},\gamma_{\mathrm{inv}}]$ with $\gamma_{\mathrm{inv}}$ being the inverse natural decay rate of the time-inverted classical dynamics [38, 45, 46, 20].
 In the semiclassical limit they are conjectured to accumulate around the typical decay rate $\gamma_{\mathrm{typ}}$ [38].
@@
 The imaginary part of $k$ is proportional to the dimensionless decay rate $\gamma$ [24, 17].
-Left resonance states $\psi^{\mathrm{L}}(\boldsymbol{r})$ fulfill
-$\psi^{\mathrm{L}}(\boldsymbol{r})\Delta=-n^{2}(\boldsymbol{r})k^{2}\psi^{\mathrm{L}}(\boldsymbol{r})$ (14)
-with incoming boundary condition.
+Left resonance states $\psi^{\mathrm{L}}(\boldsymbol{r})$ fulfill an incoming boundary condition and
+$\Delta\overline{\psi^{\mathrm{L}}(\boldsymbol{r})}=-n^{2}(\boldsymbol{r})\,k^{2}\,\overline{\psi^{\mathrm{L}}(\boldsymbol{r})}\,.$ (14)
 The LR-Husimi representation for time-continuous scattering systems is obtained by using the Husimi representation for right and left resonance states, see Eq.
@@
 For dielectric cavities the Husimi representation for right and left resonance states is defined in Ref.
-[52], specifically for right (left) resonance states we use the incoming (emanating) Husimi representation $\mathcal{H}_{n}^{\text{inc(em)}}$.
+[52].
+Specifically, for right and left resonance states we use the incoming Husimi representation $\mathcal{H}_{n}^{\text{inc}}$ inside the cavity.
+Alternatively, one could use for both states the emanating Husimi representation, giving the same LR-Husimi representation in the semiclassical limit.
+In this context the terms incoming and emanating refer to the definition of the boundary Husimi representation inside the cavity in Ref.
+[52] and are not related to the incoming and outgoing boundary conditions of left and right resonance states.
 For the three-disk scattering system the Husimi representation is defined in Ref.
@@
 [18].
-[…]
+In all figures, for the most semiclassical resonance states we use the following ranges of wavenumbers: (a) for the dielectric cavity $\mathrm{Re}\,kR_{\mathrm{cav}}\in[5\,000,5\,025]$, (b) for the three-disk scattering system $\mathrm{Re}\,ka\in[100\,000,100\,200]$ (except for $\gamma=1.8$ in Figs.
+6, 8, and 10, where $\mathrm{Re}\,ka\in[50\,000,50\,500]$ is used), and (c, d) for the standard map with partial and full escape all even $N\in[19\,996,20\,004]$.
+We expect that our results hold for higher-dimensional billiards and maps with fully chaotic phase spaces just as well, since all our arguments are independent of the number of dimensions of phase space.
+However, those systems are harder to visualize and numerically more challenging, in particular they would not allow us to go as far into the semiclassical limit.
+Therefore, we focus on the lowest dimensional billiards and maps with chaotic dynamics.
 II.4 Visualization of LR-Husimi representation
@@
 10, (i) we first insert the identity $I=U_{\text{cl}}U_{\text{cl}}^{-1}$ twice, leading to
-$\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle$ =$\stackrel{{\scriptstyle(\text{i})}}{{=}}\langle\alpha(\boldsymbol{x})|U_{\text{cl}}U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle$
+$\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle$ $\stackrel{{\scriptstyle(\text{i})}}{{=}}\langle\alpha(\boldsymbol{x})|U_{\text{cl}}U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle$
 $\stackrel{{\scriptstyle(\text{ii})}}{{\approx}}\langle\alpha(T^{-1}(\boldsymbol{x}))|U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
-=$\stackrel{{\scriptstyle(\text{iii})}}{{=}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
+$\stackrel{{\scriptstyle(\text{iii})}}{{=}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
 $\stackrel{{\scriptstyle(\text{iv})}}{{\approx}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{n}\rangle\langle L_{n}|\alpha(T^{-1}(\boldsymbol{x}))\rangle\,.$ (17)
@@
 21 and 22 to find
-$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x})\eta^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}$
+$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x})\eta^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}$ (23)
 We now use from the factorization conjecture above that the right fluctuations $\eta^{\mathrm{R}}(\boldsymbol{x})$ are statistically independent of their classical density $\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})$.
@@
 $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})=\int_{0}^{\infty}\text{d}\eta^{\mathrm{L}}\int_{0}^{\infty}\text{d}\eta^{\mathrm{R}}\,P(\eta^{\mathrm{L}})\,P(\eta^{\mathrm{R}})\;\;\delta\left(\eta^{\mathrm{LR}}-\frac{4}{\pi}\sqrt{\eta^{\mathrm{L}}\eta^{\mathrm{R}}}\right)\,.$ (29)
-For one integration the $\delta$-function can be exploited, while the other gives an integral representation of $K_{0}$, the 00-th order modified Bessel function of the second kind [56], leading to the universal distribution of LR-fluctuations,
+For one integration the $\delta$-function can be exploited, while the other gives an integral representation of $K_{0}$, the 0-th order modified Bessel function of the second kind [56], leading to the universal distribution of LR-fluctuations,
 $P^{\mathrm{LR}}(\eta^{\mathrm{LR}})=\frac{\pi^{2}}{4}\,\eta^{\mathrm{LR}}\,K_{0}\left(\frac{\pi}{2}\eta^{\mathrm{LR}}\right)\,.$ (30)
@@
 (d) Standard map with full escape for $\gamma\in\{0.25\,(\gamma_{\mathrm{nat}}),0.35,0.5,0.75,1.0\}$.
-Symbols $\circ$, $+$, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.
+Symbols $\circ$, ++, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.
 IV.4 Construction of LR-measure
@@
 III do not depend on the type of dynamics and thus seem to suggest that quantum invariance holds more generally for resonance states with a well defined semiclassical limit.
+An important open question is how the presented results can be observed in experimental scattering systems.
+As we are in a regime of strongly overlapping resonances, the properties of an individual resonance are experimentally not accessible.
+However, the average over many resonances with similar decay rates also depends on the decay rate and might be observable by the time evolution of wave packets.
+This is experimentally related to pulse propagation deduced from transmission spectra [62].
+Data availability statement
+The data that support the findings of this study are openly available at the following URL/DOI: https://doi.org/10.25532/OPARA-989.
 Acknowledgements.